rssLink RSS for all categories
 
icon_red
icon_green
icon_green
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_red
icon_green
icon_green
icon_orange
icon_green
icon_red
icon_green
icon_blue
icon_red
icon_orange
icon_green
icon_blue
icon_green
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_red
 

FS#45557 — Fiber out of service

Attached to Project— Network and racks
Incident
Entire OVH Network
CLOSED
100%
FR :

Bonjour,
Nous avons un dysfonctionnement de fibre d'un routeur entre bhs1 et chi et , nous cherchons si c'est physique ou un problème software.
Nous investiguons.

EN :

Hello,
We have some difficulties fiber on router between bhs1 and chi we are looking for a physical problem or a software problem.
We're investigating.
Date:  Friday, 10 July 2020, 11:03AM
Reason for closing:  Done
Comment by OVH - Wednesday, 08 July 2020, 17:51PM

We have an incident on the OPTICAL Network that affects 3 pairs of darkfiber

BHS<>NWK circuits are on the protection path
BHS<>CHI direct circuits are down, traffic to Chicago is going through Newark and Ashburn
BHS<>TOR direct circuits are down, traffic to Toronto is going through Newark
BHS<>MTL circuits are on the protection path.


Comment by OVH - Wednesday, 08 July 2020, 22:36PM

We currently have a fiber cut on this circuit.
1 - pair is down

2 - other pairs are affecting too on the same conduit.

Check before repair...


Comment by OVH - Wednesday, 08 July 2020, 23:56PM

Fibercut has been confirmed at 7km from BHS datacenter.

Fiber team is working to repair the damaged cable.

ETA for now is 21h00 Montreal Time.


Comment by OVH - Thursday, 09 July 2020, 05:11AM

the root cause of the fibercuts is a fire that burned the conduit. Our provider is on site and wait for the go from the power company to access the conduit and run a new cable.

Final ETR is still to be provided.


Comment by OVH - Thursday, 09 July 2020, 18:49PM

Repair always ongoing.
ETR is 14h Montreal Time


Comment by OVH - Thursday, 09 July 2020, 23:19PM

First pair is now up.
BHS<>CHI direct circuits UP
BHS<>TOR direct circuits UP

Repair always ongoing


Comment by OVH - Friday, 10 July 2020, 11:02AM

The 3 fibers are now up since 7:39 GMT+2. Fibercut cut was due to a fire at 7km from our datacenter in Beauharnois.